### v2

# Competition of $\chi_{c}(2P)$ quarkonia and continuum in $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$

Abstract
We discuss the production of $D\bar{D}$ pairs in $e^{+}e^{-}$ collisions, where $D$ refers to either $D^{0}$ or $D^{+}$. The continuum mechanism with the $t/u$-channel vector-meson $D^{*}$ exchanges are considered. The results of the calculation depend on the parameter of the off-shell form-factor for the virtual $D^{*}$ mesons. The $D^{*}D\gamma$ coupling constants are found from the $D^{*}\to D\gamma$ decays. We find relatively large contribution for the $D^{0}{\bar{D}}^{0}$ channel and much smaller contribution in the $D^{+}D^{-}$ channel. In the second case we consider also the $D^{\pm}$ exchanges. We conclude that the bump at $M_{D^{0}{\bar{D}}^{0}}=3.8$ GeV observed by the Belle and BaBar Collaborations has rather continuum origin than it corresponds to the broad resonance $\chi_{c0}(3860)$. We discuss also production of the $\chi_{c2}(3930)$ resonance which is a candidate for the $\chi_{c2}(2P)$ state. This state can decay into both $D\bar{D}$ channels, however the branching fractions are not well known at present. From a comparison of our model results to the BaBar data we find $B(\chi_{c2}(3930)\to D\bar{D})=0.58\pm 0.13$ using the two-photon width $\Gamma_{\gamma\gamma}=0.544$ keV (obtained for the Buchmüller-Tye potential) evaluated within the light-front approach (NRQCD limit). Our finding of $\Gamma_{\gamma\gamma}\times B(\chi_{c2}(3930)\to D\bar{D})=0.32\pm 0.07$ keV is close to the Belle and BaBar results. Realistic predictions of the differential distributions in several variables and integrated cross-sections are given for the Belle II kinematics.
I Introduction
The ground state properties of $P$-wave charmonia $\chi_{cJ}$ ($J=0,1,2$) are relatively well known. Their decay widths and branching fractions were measured [1]. In contrast, the excited $P$-wave charmonia (their masses, widths, and branching fractions) are poorly known. Recently, the production of axial $\chi_{c1}(2P)$ state in $pp$ [2] and $e^{+}e^{-}$ [3] collisions was discussed. These studies indicate that also for the famous $X(3872)$, although potentially an exotic state, the $c\bar{c}$ component can play an important role. This naturally raises the question of whether similar situation can occur for the $J^{P}=0^{+}$ and $2^{+}$ states. A number of observed states are considered as candidates for the first excited $P$-wave charmonia $\chi_{c0,2}(2P)$. A candidate for the $\chi_{c2}(2P)$ state ($\chi_{c2}(3930)$ in PDG [1]) was observed by the Belle Collaboration [4] and later confirmed by the BaBar Collaboration [5] in the two-photon fusion process $e^{+}e^{-}\to e^{+}e^{-}(\gamma\gamma\to D\bar{D})$ as a resonance in the $D^{+}D^{-}$ and $D^{0}\bar{D}^{0}$ invariant mass distributions. In the $D^{0}{\bar{D}}^{0}$ channel, one can also observe a broad enhancement at $M\sim 3.8$ GeV. The intriguing problem of searching for $\chi_{c0}(2P)$ was a topic of several works; see [6, 7]. The candidate for the $\chi_{c0}(2P)$ state is either the $X(3860)$ ($\chi_{c0}(3860)$ in PDG) or the $X(3915)$ ($\chi_{c0}(3915)$ in PDG). The $X(3915)$ state is relatively narrow ($\Gamma\simeq 20$ MeV) and was observed in two-photon fusion in the $\omega J/\psi$ decay mode [8] as well as in the $B$ meson decays, $B\to KJ/\psi\omega$ (Belle [9], BaBar [10]) and $B^{+}\to D^{+}D^{-}K^{+}$ (LHCb [11]). The observation of a broad resonance identified as a new charmonium-like state $X(3860)$ with a preference for $J^{PC}=0^{++}$ over $2^{++}$ was reported by the Belle Collaboration [12] in the process $e^{+}e^{-}\to J/\psi D\bar{D}$. In line with the arguments presented there, the $X(3860)$ state seems to be a better candidate for the $\chi_{c0}(2P)$ charmonium state than the $X(3915)$. There is still much uncertainty about the above assignments for the $\chi_{c0,2}(2P)$ states due to limited information on their production and decay properties, and therefore further confirmation is required.
The nature of the $\chi_{c0}(3915)$ and $\chi_{c2}(3930)$ charmonium-like states is also not well understood. In Ref. [13] the authors suggest that the two resonances $X(3915)$ and $X(3930)$ can be regarded as the same $J^{PC}=2^{++}$ state. It was shown in [14], using the framework of a constituent quark model, that this may be related to the molecular structure for these two states. Further considerations on the $X(3915)/X(3930)$ puzzle can be found in [15] and in the references therein. Prompt hadroproduction of the $\chi_{c2}(3930)$ resonance was observed by the LHCb Collaboration in the decay modes $\chi_{c2}(3930)\to D^{+}D^{-}$ and $D^{0}\bar{D}^{0}$ in proton-proton collisions [16]. A more recent LHCb amplitude analysis of the process $B^{+}\to D^{+}D^{-}K^{+}$ finds distinct $J^{PC}=0^{++}$ and $2^{++}$ states decaying to $D^{+}D^{-}$ [11], which can be identified as the $\chi_{c0}(3915)$ and $\chi_{c2}(3930)$ charmonium states, respectively.
Regarding the reactions $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$, where $D^{0}, D+D^{+}$, we are not convinced of the existence of a broad resonance ($\Gamma\sim 200$ MeV) at $M\sim 3.85$ GeV reported by the Belle and BaBar Collaborations. The reason being that while an enhancement is observed in the $D^{0}{\bar{D}}^{0}$ invariant mass distribution, the effect is not clearly seen in the $D^{+}D^{-}$ channel. In order to increase the statistics several groups combined the $D^{0}{\bar{D}}^{0}$ and $D^{+}D^{-}$ final states and studied the combined data. For example, some of the first descriptions of the combined Belle and BaBar data on $\gamma\gamma\to D\bar{D}$ reactions were reported in [6, 17, 13]. More recently, the data-driven analysis of $\gamma\gamma\to D\bar{D}$ performed in a partial-wave dispersive formalism [18], indicates no evidence for the broad resonance $X^{*}(3860)$, but rather reports a bound state below the $D\bar{D}$ threshold. In [19] a possible bound state of $D\bar{D}$ generated by the meson-meson interaction is discussed. In Ref. [20] the available data in the $D\bar{D}$ and $D_{s}^{+}D_{s}^{-}$ channels from both $\gamma\gamma$-fusion reaction and $B^{+}\to D^{+}D^{-}K^{+}$, $B^{+}\to D_{s}^{+}D_{s}^{-}K^{+}$ decays were analyzed considering the $D\bar{D}$-$_{s}\bar{D}_{s}-DD$$^{*}\bar{D}^{*}$-$D_{s}^{*}\bar{D}_{s}^{*}$ coupled channels with scalar and tensor states.
In this paper, we consider two different mechanisms contributing to the $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$ reactions. The first mechanism is the so-called continuum contribution described by $t$- and $u$-channel meson exchanges and contact terms. For the $D^{0}{\bar{D}}^{0}$ final state we consider the vector-meson $D^{*}(2007)^{0}$ $t/u$-channel exchanges as the underlying continuum process while in the $D^{+}D^{-}$ case we consider the pseudoscalar $D^{\pm}$ $t/u$-channel exchanges plus a contact term. The second production mechanism proceeds through an intermediate resonance state. Here, we shall examine only the $\chi_{c0}(3860)$ and $\chi_{c2}(3930)$ resonances, assuming that they correspond to $\chi_{c0}(2P)$ and $\chi_{c2}(2P)$, respectively. We stress that we aim to describe the Belle and BaBar data simultaneously, for both the neutral and charged channels separately.
In this paper we discuss the production of $D\bar{D}$ meson pairs in $e^{+}e^{-}$ collisions, concentrating on low $\gamma\gamma$ energies ($W_{\gamma\gamma}<4.3$ GeV). In [21], the interested reader may find predictions for the exclusive $D\bar{D}$ production in ultraperipheral ultrarelativistic heavy-ion collisions (UPCs). There, the $\gamma\gamma\to D\bar{D}$ subprocesses were estimated for $W_{\gamma\gamma}>4$ GeV in the heavy-quark approximation and in the Brodsky-Lepage formalism. It is also worth noting that the production of open-charm meson pairs in UPCs is one of the potential opportunities to search for the lightest possible exotic charmonium state; see [22].
II Production of $D\bar{D}$ pairs in $e^{+}e^{-}$ collisions via $\gamma\gamma$ fusion
Figure 1: Diagrams representing $D^{+}D^{-}$ production in $e^{+}e^{-}$ collisions. For the process $\gamma\gamma\to D^{+}D^{-}$ we consider the $t$- and $u$-channel $D^{\pm}$ and $D^{*}(2010)^{\pm}$ exchanges, the contact term, and the exchange of $\chi_{c0,2}$ mesons in the $s$ channel. The diagrams for the $D^{0}{\bar{D}}^{0}$ production are similar, but in the case of non-resonant production we consider only the $D^{*}(2007)^{0}$ $t/u$-channel exchanges.
In the present studies, we perform calculations for the considered exclusive $2\to 4$ reaction; see the diagrams shown in Fig. 1. The fully differential cross section reads
$d\sigma=\frac{1}{2\sqrt{s(s-4m_{e}^{2})}}\overline{|{\cal M}_{2\to 4}|^{2}}(2\pi)^{4}\delta^{(4)}(p_{1}+p_{2}+p_{3}+p_{4}-p_{a}-p_{b})\prod_{i=1}^{4}\frac{\text{d}^{3}p_{i}}{(2\pi)^{3}2p_{i}^{0}}\,.$ (2.1)
Here four-momenta of incoming and outgoing particles are denoted by $p_{a},p_{b},p_{1},\ldots,p_{4}$, the energy variable is $s=(p_{a}+p_{b})^{2}$, $m_{e}$ is the electron mass, and $\overline{|{\cal M}_{2\to 4}|^{2}}$ is the amplitude squared averaged over initial and summed over final lepton polarization states. To calculate the total cross section one has to perform an eight dimensional phase space integration as is done for instance for the $pp\to pp\pi^{+}\pi^{-}$ reaction [23] and for the $pp\to pp\mu^{+}\mu^{-}$ reaction [24]. In the present paper, we use the following integration variables: the transverse momenta of outgoing leptons ($p_{t1}$, $p_{t2}$), the azimuthal angles of outgoing leptons ($\phi_{1}$, $\phi_{2}$), the rapidities of the mesons ($y_{3}$, $y_{4}$), the length and the azimuthal angle of the difference of the transverse momenta of the mesons ($p_{mt}$, $\phi_{m}$).
We are interested in the photon-photon fusion production mechanism relevant for small–angle scattering of leptons. The generic $2\to 4$ amplitude has the form
${\cal M}_{\lambda_{a}\lambda_{b}\to\lambda_{1}\lambda_{2}D\bar{D}}=ij^{(1)}_{\alpha}(\lambda_{1},\lambda_{a})\,\frac{-ig^{\alpha\mu}}{q_{1}^{2}}\,{\cal M}_{\mu\nu}(\gamma^{*}\gamma^{*}\to D\bar{D})\,\frac{-ig^{\nu\beta}}{q_{2}^{2}}\,ij^{(2)}_{\beta}(\lambda_{2},\lambda_{b})\,$, (2.2)
where $\lambda_{a},\lambda_{b},\lambda_{1},\lambda_{2}\in\{1/2,-1/2\}$ are the lepton polarizations, and $q_{1}=p_{a}-p_{1}$, $q_{2}=p_{b}-p_{2}$. Below, we also use the Mandelstam variables $t_{1}=q_{1}^{2}$, $t_{2}=q_{2}^{2}$, and $p_{34}=q_{1}+q_{2}=p_{3}+p_{4}$. Furthermore,
$j^{(1)}_{\alpha}(\lambda_{1},\lambda_{a})$ $=$ $\bar{v}(p_{1},\lambda_{1})\gamma_{\alpha}v(p_{a},\lambda_{a})\approx(p_{1}+p_{a})_{\alpha}\,\delta_{\lambda_{1}\lambda_{a}}\,$,
$j^{(2)}_{\beta}(\lambda_{2},\lambda_{b})$ $=$ $\bar{u}(p_{2},\lambda_{2})\gamma_{\beta}u(p_{b},\lambda_{b})\approx(p_{2}+p_{b})_{\beta}\,\delta_{\lambda_{2}\lambda_{b}}\,$, (2.3)
are the lepton currents. Let us now collect the amplitudes for the various $\gamma^{*}\gamma^{*}\to D\bar{D}$ subprocesses considered by us.
II.1 Resonant production: $\gamma^{*}\gamma^{*}\to\chi_{c2}\to D\bar{D}$
The amplitude for the $\gamma^{*}(q_{1})\gamma^{*}(q_{2})\to\chi_{c2}\to D(p_{3})\bar{D}(p_{4})$ subprocess is given by
${\cal M}_{\mu\nu}(\gamma^{*}\gamma^{*}\to\chi_{c2}\to D\bar{D})$ $=$ $i\Gamma^{(\gamma^{*}\gamma^{*}\to\chi_{c2})}_{\mu\nu\rho\sigma}(q_{1},q_{2})\;i\Delta^{(\chi_{c2})\,\rho\sigma,\alpha\beta}(p_{34})\;i\Gamma^{(\chi_{c2}\to D\bar{D})}_{\alpha\beta}(p_{3},p_{4})\,.$ (2.4)
Here, we have introduced the propagator for a narrow tensor-meson resonance in the relativistic Breit-Wigner form
$i\Delta_{\mu\nu,\kappa\lambda}^{(\chi_{c2})}(p_{34})$ $=$ $\frac{i}{p_{34}^{2}-M_{\chi_{c2}}^{2}+iM_{\chi_{c2}}\Gamma_{\chi_{c2}}}\left[\frac{1}{2}(\hat{g}_{\mu\kappa}\hat{g}_{\nu\lambda}+\hat{g}_{\mu\lambda}\hat{g}_{\nu\kappa})-\frac{1}{3}\hat{g}_{\mu\nu}\hat{g}_{\kappa\lambda}\right]\,$, (2.5)
where $\hat{g}_{\mu\nu}=-g_{\mu\nu}+p_{34\mu}p_{34\nu}/p_{34}^{2}$. The total decay width of the $\chi_{c2}(3930)$ resonance and its mass is taken from the PDG [1] as:
$M_{\chi_{c2}}$ $=$ $3922.5\pm 1.0~{\rm MeV}\,$,
$\Gamma_{\chi_{c2}}$ $=$ $35.2\pm 2.2~{\rm MeV}\,.$ (2.6)
II.1.1 Two photons vertex: $\gamma^{*}\gamma^{*}\to\chi_{c2}$
The production of the $\chi_{c2}$ resonance is encoded in the $\gamma^{*}\gamma^{*}\to\chi_{c2}$ vertex, which in general depends on five invariant form factors that are functions of the photon virtualities $q_{1}^{2},q_{2}^{2}$. For the corresponding Lorentz structures, see e.g. [25, 26, 27]. Here we concentrate on the resonance $\chi_{c2}(3930)$ which can be identified as the $\chi_{c2}(2P)$ charmonium $c\bar{c}$ state. We anticipate that the main contribution to the cross section will come from quasireal transversally polarized photons. In addition, we expect that the helicity-2 component will dominate over the helicity-0 one. For example in the nonrelativistic quark model, the contribution of the $J_{z}=0$ polarization of the heavy tensor meson vanishes, see e.g. [28]. We therefore limit ourselves only to helicity-2 component, so that the vertex can be written as
$\Gamma_{\mu\nu\kappa\lambda}^{(\gamma^{*}\gamma^{*}\to\chi_{c2})}(q_{1},q_{2})=e^{2}\frac{1}{2}\Big(R_{\mu\kappa}R_{\nu\lambda}+R_{\nu\kappa}R_{\mu\lambda}-R_{\mu\nu}R_{\kappa\lambda}\Big)F_{TT,2}(Q_{1}^{2},Q_{2}^{2})\,$, (2.7)
where $e^{2}=4\pi\alpha_{\rm em}$, the photon virtualities $Q_{1,2}^{2}\equiv-q_{1,2}^{2}\geqslant 0$, and
$R_{\mu\nu}\equiv R_{\mu\nu}(q_{1},q_{2})=-g_{\mu\nu}+\frac{1}{X}\Big((q_{1}\cdot q_{2})(q_{1\mu}q_{2\nu}+q_{1\nu}q_{2\mu})-q_{1}^{2}q_{2\mu}q_{2\nu}-q_{2}^{2}q_{1\mu}q_{1\nu}\Big)\,$, (2.8)
$X=(q_{1}\cdot q_{2})^{2}-q_{1}^{2}q_{2}^{2}$. This vertex function satisfies the relations
$\Gamma_{\mu\nu\kappa\lambda}(q_{1},q_{2})=\Gamma_{\nu\mu\kappa\lambda}(q_{2},q_{1})\,$, (2.9)
$q_{1}^{\mu}\Gamma_{\mu\nu\kappa\lambda}(q_{1},q_{2})=0\,$, (2.10)
$q_{2}^{\nu}\Gamma_{\mu\nu\kappa\lambda}(q_{1},q_{2})=0\,$, (2.11)
$\Gamma_{\mu\nu\kappa\lambda}(q_{1},q_{2})g^{\kappa\lambda}=0\,$, (2.12)
$\Gamma_{\mu\nu\kappa\lambda}(q_{1},q_{2})(q_{1}+q_{2})^{\kappa}(q_{1}+q_{2})^{\lambda}=0\,$, (2.13)
and is also symmetric when exchanging $\kappa\leftrightarrow\lambda$.
The transition form factor will be parametrized as
$F_{TT,2}(Q_{1}^{2},Q_{2}^{2})=F_{TT,2}(0,0)\frac{\Lambda^{2}}{Q_{1}^{2}+Q_{2}^{2}+\Lambda^{2}}\,.$ (2.14)
In the calculations we take $\Lambda=M_{\chi_{c2}(3930)}=3.9225$ GeV which, in principle, is a free parameter. We would like to mention that $F_{TT,2}(0,0)$, the form factor at the on-shell point, is not dimensionless.
Assuming that the tensor $\chi_{c2}(2P)$ state is quarkonium-like (pure $c\bar{c}$ state) one can calculate $F_{TT,2}(0,0)$ in the NRQCD limit [see (4.2) of [29]]
$F_{TT,2}(0,0)=8e_{f}^{2}\sqrt{\frac{3N_{c}}{\pi M_{\chi_{c2}}^{3}}}|R_{2P}^{\prime}(0)|$ (2.15)
with $e_{f}=2/3$ and $N_{c}=3$. The first derivative of the radial wave function at the origin is estimated for several $c\bar{c}$ interaction potential models from the literature, listed in Table 1 below. Following the approach in [27], we solve Schrödinger equation, but for $2P$ state, and get $|R_{2P}^{\prime}(0)|$ values. As shown in Table 1, there is a considerable spread of results for the different potential models.
Table 1: Results of the derivative of the wave function at the origin $|R_{2P}^{\prime}(0)|$ for $\chi_{c2}(2P)$ state obtained from the light front quark model for five different potential model, the corresponding values of transition form factor (2.15), and of two-photon width for the helicity-2 contribution in NRQCD limit.
Potential type $|R^{\prime}_{2P}(0)|\,(\rm GeV^{5/2})$ $|F_{TT,2}(0,0)|\,(\rm GeV)$ $\Gamma_{\gamma\gamma}(\lambda=\pm 2)\,(\rm keV)$
Cornell 0.405 0.314 0.840
Buchmüller-Tye 0.326 0.253 0.544
harmonic oscillator 0.275 0.213 0.387
logarithmic 0.263 0.204 0.354
power like 0.228 0.177 0.266
In the NRQCD limit, the helicity-0 term vanishes and the helicity-2 is the dominant one. Then, $F_{TT,2}(0,0)$ is related to the two-photon decay width of the $\chi_{c2}$ state
$\Gamma(\chi_{c2}(2P)\to\gamma\gamma)=(4\pi\alpha_{\rm em})^{2}\frac{|F_{TT,2}(0,0)|^{2}}{80\pi M_{\chi_{c2}}}\,;$ (2.16)
see the discussion in Sec. 4 of [29]. Using $|R_{2P}^{\prime}(0)|=(0.228-0.405)$ GeV5/2 we find $\Gamma_{\gamma\gamma}(\lambda=\pm 2)=(0.27-0.84)$ keV assuming the dominance of the helicity-2 term. Here, the Cornell potential gives the largest value of the two-photon width. It is worth noting that the BLFQ prediction [30] for the two-photon width of $\chi_{c2}(3930)$, $\Gamma_{\gamma\gamma}=0.58(25)$ keV (see Table I of [30]), is most compatible with our result obtained for the Buchmüller-Tye potential.
For comparison, a value $|R_{2P}^{\prime}(0)|^{2}=0.1767$ GeV5 was found in [31] computed for the mass $M=3.9554$ GeV with $m_{c}=1.84$ GeV within the Cornell potential. This value, $|R_{2P}^{\prime}(0)|\simeq 0.420$ GeV5/2, appears to be larger than our findings. Nevertheless, it was used for the inclusive $\chi_{c1}(3872)$ production [2] leading to a good description of the proton-proton data within the uncertainty of the model (different gluon uPDFs).
Alternatively, for the $\gamma^{*}\gamma^{*}\to\chi_{c2}$ vertex one can use the form given in (C25) of [26]. In this case, the helicity-2 component reads
$\Gamma_{\mu\nu\kappa\lambda}^{\rm PPV}(q_{1},q_{2})=e^{2}\Big(R_{\mu\kappa}(q_{1},q_{2})R_{\nu\lambda}(q_{1},q_{2})+\frac{\hat{s}}{8X}R_{\mu\nu}(q_{1},q_{2})(q_{1}-q_{2})_{\kappa}(q_{1}-q_{2})_{\lambda}\Big)\frac{q_{1}\cdot q_{2}}{M_{\chi_{c2}}}\tilde{F}_{TT,2}(Q_{1}^{2},Q_{2}^{2})\,$, (2.17)
where $\hat{s}=(q_{1}+q_{2})^{2}=2(q_{1}\cdot q_{2})+q_{1}^{2}+q_{2}^{2}$. We have $q_{1}\cdot q_{2}=(M_{\chi_{c2}}^{2}+Q_{1}^{2}+Q_{2}^{2})/2$. Introducing an extra factor $(q_{1}\cdot q_{2})/M_{\chi_{c2}}$ the transition form factor $\tilde{F}_{TT,2}(Q_{1}^{2},Q_{2}^{2})$ remains dimensionless. For the vertex function (2.17) we find the relations
$\Gamma_{\mu\nu\kappa\lambda}^{\rm PPV}(q_{1},q_{2})g^{\kappa\lambda}\neq 0\,$, (2.18)
$\Gamma_{\mu\nu\kappa\lambda}^{\rm PPV}(q_{1},q_{2})(q_{1}+q_{2})^{\kappa}(q_{1}+q_{2})^{\lambda}\neq 0\,.$ (2.19)
In order to get compatibility with the formula given by Eq. (2.7), we write
$\tilde{F}_{TT,2}(Q_{1}^{2},Q_{2}^{2})=\frac{M_{\chi_{c2}}}{q_{1}\cdot q_{2}}F_{TT,2}(Q_{1}^{2},Q_{2}^{2})\,.$ (2.20)
The vertices (2.7) and (2.17) are written for ‘on-shell’ $\chi_{c2}$ meson, that is for $p_{34}^{2}=M_{\chi_{c2}}^{2}$. For $p_{34}^{2}\neq M_{\chi_{c2}}^{2}$ the form factor $F^{(\chi_{c2})}(p_{34}^{2})$ can be inserted in addition. The form factor should be normalized to $F^{(\chi_{c2})}(M_{\chi_{c2}}^{2})=1$. In our calculation we put $F^{(\chi_{c2})}(p_{34}^{2})=1$.
II.1.2 $\chi_{c2}\to D\bar{D}$ vertex
The $\chi_{c2}D\bar{D}$ vertex (see Eq. (3.37) of [32] for the analogous $f_{2}\pi\pi$ vertex) can be written as
$i\Gamma_{\mu\nu}^{(\chi_{c2}\to D\bar{D})}(p_{3},p_{4})=-i\,\frac{g_{\chi_{c2}D\bar{D}}}{2M_{0}}\left[(p_{3}-p_{4})_{\mu}(p_{3}-p_{4})_{\nu}-\frac{1}{4}g_{\mu\nu}(p_{3}-p_{4})^{2}\right]F^{(\chi_{c2})}(p_{34}^{2})$ (2.21)
with $M_{0}\equiv 1$ GeV. We have $\Gamma_{\mu\nu}^{(\chi_{c2}\to D\bar{D})}(p_{3},p_{4})g^{\mu\nu}=0$. Here we can assume the form factor for the off-shell $\chi_{c2}$ meson. In our calculation we put $F^{(\chi_{c2})}(p_{34}^{2})=1$.
The decay width of $\chi_{c2}\to D^{+}D^{-}$ is given by
$\Gamma(\chi_{c2}\to D^{+}D^{-})=\frac{M_{\chi_{c2}}}{480\pi}\,|g_{\chi_{c2}D\bar{D}}|^{2}\left(\frac{M_{\chi_{c2}}}{M_{0}}\right)^{2}\left(1-\frac{4m_{D}^{2}}{M_{\chi_{c2}}^{2}}\right)^{5/2}\,.$ (2.22)
We take $m_{D}=m_{D^{+}}$. The $\chi_{c2}(3930)$ has isospin $I=0$. Assuming isospin invariance in the decay, we get
$\Gamma(\chi_{c2}\to D^{+}D^{-})=\Gamma(\chi_{c2}\to D^{0}\bar{D}^{0})=\frac{1}{2}\Gamma(\chi_{c2}\to D\bar{D})\,.$ (2.23)
We noted that the most significant partial widths for $\chi_{c2}(2P)$ are $\Gamma(\chi_{c2}\to D\bar{D})$ and $\Gamma(\chi_{c2}\to D\bar{D}^{*})$. Different calculations in the literature support the fact that the total width of $\chi_{c2}(2P)$ is largely saturated by the $D\bar{D}$ mode. For instance, in Eq. (24) and Fig. 24 of [33] one can find $\Gamma(\chi_{c2}\to D\bar{D})\approx 26(12)$ MeV and $\Gamma(\chi_{c2}\to D\bar{D}^{*})\approx 22(14)$ MeV. In Table I of [34] one can find $\Gamma(\chi_{c2}\to D\bar{D})=24$ MeV and $\Gamma(\chi_{c2}\to D\bar{D}^{*})=13.6-16.6$ MeV (depending on model). In Table IV of [35], the following values corresponding to $M_{\chi_{c2}(2P)}=3930$ MeV are stated $\Gamma(\chi_{c2}\to D\bar{D})=20.4$ MeV and $\Gamma(\chi_{c2}\to D\bar{D}^{*})=6.9$ MeV. On the other hand, the authors of [14, 37] found a different decomposition of partial widths.
Thus, for the estimate, we take $\Gamma(\chi_{c2}\to D\bar{D})=14-26$ MeV, and using (2.6), (2.22), and (2.23) we obtain $|g_{\chi_{c2}D\bar{D}}|=8.34-11.37$. With $\Gamma_{\chi_{c2}}=35.2$ MeV (3.2) we get the branching fraction in the range $B(\chi_{c2}\to D\bar{D})=0.40-0.74$.
II.2 Resonant production: $\gamma^{*}\gamma^{*}\to\chi_{c0}\to D\bar{D}$
Here we write formulas for the production of the scalar $\chi_{c0}(2P)$ state assumed by us to have the mass of $\chi_{c0}(3860)$. The total decay width of the $\chi_{c0}(3860)$ resonance and its mass given in the PDG [1] is:
$M_{\chi_{c0}}$ $=$ $3862^{+50}_{-35}~{\rm MeV}\,$,
$\Gamma_{\chi_{c0}}$ $=$ $200^{+180}_{-110}~{\rm MeV}\,.$ (2.24)
In the calculations, we also consider another scenario:
$M_{\chi_{c0}}$ $=$ $3862~{\rm MeV}\,$,
$\Gamma_{\chi_{c0}}$ $=$ $50~{\rm MeV}\,.$ (2.25)
The narrower total width of $\chi_{c0}(3860)$ seems to be consistent with BaBar data for $D^{+}D^{-}$ channel; see Fig. 3(b) below. Similar value is given in Table VI of [36] (see $2^{3}P_{0}$ state).
The amplitude for the $\gamma^{*}(q_{1})\gamma^{*}(q_{2})\to\chi_{c0}\to D(p_{3})\bar{D}(p_{4})$ subprocess is given by
${\cal M}_{\mu\nu}(\gamma^{*}\gamma^{*}\to\chi_{c0}\to D\bar{D})=i\Gamma^{(\gamma^{*}\gamma^{*}\to\chi_{c0})}_{\mu\nu}(q_{1},q_{2})\,i\Delta^{(\chi_{c0})}(p_{34})\,i\Gamma^{(\chi_{c0}\to D\bar{D})}(p_{3},p_{4})\,.$ (2.26)
The scalar-meson propagator is parametrized as
$i\Delta^{(\chi_{c0})}(p_{34})=\frac{i}{p_{34}^{2}-M_{\chi_{c0}}^{2}+iM_{\chi_{c0}}\Gamma_{\chi_{c0}}}$ (2.27)
with a constant decay width.
The $\gamma^{*}\gamma^{*}\to\chi_{c0}$ vertex, limiting only to the transverse component, can be written as
$\Gamma_{\mu\nu}^{(\gamma^{*}\gamma^{*}\to\chi_{c0})}(q_{1},q_{2})=-e^{2}R_{\mu\nu}F_{TT}(Q_{1}^{2},Q_{2}^{2})\,.$ (2.28)
Here, $R_{\mu\nu}$ is defined by (2.8) and the form factor $F_{TT}$ is parametrized as
$F_{TT}(Q_{1}^{2},Q_{2}^{2})=F_{TT}(0,0)\frac{\Lambda^{2}}{Q_{1}^{2}+Q_{2}^{2}+\Lambda^{2}}\,,\quad\Lambda=M_{\chi_{c0}}\,.$ (2.29)
Assuming that the state is quarkonium-like one can calculate $F_{TT}(0,0)$ in the NRQCD limit as [see (4.11) of [27]]
$F_{TT}(0,0)=e_{f}^{2}\sqrt{N_{c}}\frac{12}{\sqrt{\pi}}\frac{R_{2P}^{\prime}(0)}{M_{\chi_{c0}}^{3/2}}\,.$ (2.30)
The two-photon decay width for $\chi_{c0}(2P)\to\gamma\gamma$ is given by
$\Gamma(\chi_{c0}(2P)\to\gamma\gamma)=\frac{\pi\alpha_{\rm em}^{2}}{M_{\chi_{c0}}}|F_{TT}(0,0)|^{2}\,.$ (2.31)
In Table 2 we listed our estimates for $|F_{TT}(0,0)|$ and $\Gamma_{\gamma\gamma}$ for the $\chi_{c0}(2P)$ state. These values are larger than for $\chi_{c2}(2P)$ given in Table 1. It is worth noting that the BLFQ approach [30] for the two-photon width of $\chi_{c0}(2P)$ predicts $\Gamma_{\gamma\gamma}=0.68(22)$ keV.
Table 2: Results for $\chi_{c0}(2P)$ state with $M_{\chi_{c0}}=3862$ MeV obtained from the light front quark model for different potential models in NRQCD limit.
Potential type $|R^{\prime}_{2P}(0)|\,(\rm GeV^{5/2})$ $|F_{TT}(0,0)|\,(\rm GeV)$ $\Gamma_{\gamma\gamma}\,(\rm keV)$
Cornell 0.405 0.278 3.351
Buchmüller-Tye 0.326 0.224 2.171
harmonic oscillator 0.275 0.189 1.545
logarithmic 0.263 0.181 1.413
power like 0.228 0.157 1.062
For the $\chi_{c0}D\bar{D}$ vertex we have ($M_{0}\equiv 1$ GeV)
$i\Gamma^{(\chi_{c0}\to D\bar{D})}(p_{3},p_{4})=ig_{\chi_{c0}D\bar{D}}M_{0}\,F^{(\chi_{c0})}(p_{34}^{2})\,$, (2.32)
where the coupling constant $g_{\chi_{c0}D\bar{D}}$ is related to the partial decay width of the $\chi_{c0}$ meson
$\Gamma{(\chi_{c0}\to D\bar{D})}=\frac{M_{0}^{2}}{16\pi M_{\chi_{c0}}}|g_{\chi_{c0}D\bar{D}}|^{2}\left(1-\frac{4m_{D}^{2}}{M_{\chi_{c0}}^{2}}\right)^{1/2}\,.$ (2.33)
For the isoscalar $\chi_{c0}$ we have
$\Gamma(\chi_{c0}\to D^{+}D^{-})=\Gamma(\chi_{c0}\to D^{0}\bar{D}^{0})=\frac{1}{2}\Gamma(\chi_{c0}\to D\bar{D})\,.$ (2.34)
In (2.28) and (2.32) we can assume the form factor for off-shell $\chi_{c0}$ meson. In our calculations we put $F^{(\chi_{c0})}(p_{34}^{2})=1$.
There is considerable uncertainty in the literature regarding the estimation of open-charm strong decay width. For instance, in Sec. III A and Table II of [35] one can find $\Gamma(\chi_{c0}\to D\bar{D})=21.0(16.4)$ MeV for $M_{\chi_{c0}}=3836.8(3862.0)$ MeV, which is rather narrow. Similar values for the open-charm strong decay width are predicted in [34] within the linear (screened) potential model, and the results are: $\Gamma(\chi_{c0}\to D\bar{D})=22(28)$ MeV for $M_{\chi_{c0}}=3869(3848)$ MeV. In the analysis of [37] the width is predicted to be 16.6 MeV, which is in agreement with the results of [34, 35] where the adopted decay model is also the $^{3}P_{0}$ pair creation model. In [38], a smaller width about 11 MeV was obtained, while the analysis of [39] in a $^{3}P_{0}$ model obtained a large width of 110–180 MeV [which is consistent with (2.24)]. The recent analysis of [40] predicts the strong decay width of 73 MeV for $\chi_{c0}(2P)$ state. This suggest that further precise measurement of the $\chi_{c0}(3860)$ state parameters and more theoretical studies of the $\chi_{c0}(2P)$ state are still needed.
Taking $M_{\chi_{c0}}=3862$ MeV, $\Gamma(\chi_{c0}\to D\bar{D})=16.4$ MeV, and using Eqs. (2.33), (2.34) we obtain $|g_{\chi_{c0}D\bar{D}}|=2.52$. This value for $g_{\chi_{c0}D\bar{D}}$ is used in our calculations presented in Sec. III. Considering the Buchmüller-Tye potential from Table 2 and assuming the branching fraction $B(\chi_{c0}\to D\bar{D})=\Gamma(\chi_{c0}\to D\bar{D})/\Gamma_{\chi_{c0}}=16.4~{\rm MeV}/50.0~{\rm MeV}=0.328$ we get
$\Gamma_{\gamma\gamma}\times B(\chi_{c0}\to D\bar{D})=2.171~{\rm keV}\times 0.328=0.71~{\rm keV}\,.$ (2.35)
It should be emphasized that the estimate of $\Gamma_{\gamma\gamma}\times B(\chi_{c0}\to D\bar{D})$ is highly uncertain in view of the uncertainty of $\Gamma_{\gamma\gamma}$ (dependent on the choice of potential model) and $B(\chi_{c0}\to D\bar{D})$ regarding the choice of total width $\Gamma_{\chi_{c0}}$, (2.24) or (2.25), and of $\Gamma(\chi_{c0}\to D\bar{D})$ (see the discussion above). Note that the present experimental widths are not well determined. Clearly, more investigations are necessary to confirm the property of the $\chi_{c0}(2P)$.
II.3 Continuum production of neutral $D$ mesons: $\gamma^{*}\gamma^{*}\to D^{0}\bar{D}^{0}$
The $\gamma\gamma\to D^{0}\bar{D}^{0}$ amplitude with the $D^{*0}\equiv D^{*}(2007)^{0}$ $t/u$-channel exchanges can be expressed as
${\cal M}_{\mu\nu}(\gamma^{*}\gamma^{*}\to D^{0}\bar{D}^{0})$ $=$ $i\Gamma^{(D^{*0}D^{0}\gamma)}_{\kappa_{1}\mu}(\hat{p}_{t},q_{1})\,i\tilde{\Delta}^{(D^{*0})\,\kappa_{1}\kappa_{2}}(s_{34},\hat{p}_{t}^{2})\,i\Gamma^{(D^{*0}D^{0}\gamma)}_{\kappa_{2}\nu}(-\hat{p}_{t},q_{2})$ (2.36)
$+\,i\Gamma^{(D^{*0}D^{0}\gamma)}_{\kappa_{1}\mu}(-\hat{p}_{u},q_{1})\,i\tilde{\Delta}^{(D^{*0})\,\kappa_{1}\kappa_{2}}(s_{34},\hat{p}_{u}^{2})\,i\Gamma^{(D^{*0}D^{0}\gamma)}_{\kappa_{2}\nu}(\hat{p}_{u},q_{2})\,$,
where $\hat{p}_{t}=p_{a}-p_{1}-p_{3}$, $\hat{p}_{u}=p_{4}-p_{a}+p_{1}$, and $s_{34}=p_{34}^{2}$.
The $D^{*0}D^{0}\gamma$ vertex, including a form factor, is taken as:
$i\Gamma_{\mu\nu}^{(D^{*0}D^{0}\gamma)}(\hat{p},q)=-ie\frac{g_{D^{*0}D^{0}\gamma}}{m_{D^{*0}}}\,\varepsilon_{\mu\nu\rho\sigma}\hat{p}^{\rho}q^{\sigma}F^{(D^{*0}D^{0}\gamma)}(\hat{p}^{2},q^{2})\,.$ (2.37)
We use the factorized form for the $D^{*0}D^{0}\gamma$ form factor
$F^{(D^{*0}D^{0}\gamma)}(\hat{p}^{2},q^{2})=F^{(D^{*0})}(\hat{p}^{2})F^{(\gamma)}(q^{2})$ (2.38)
with $F^{(D^{*0}D\gamma)}(m_{D^{*0}}^{2},0)=1$. We take
$F^{(D^{*0})}(\hat{p}^{2})$ $=$ $\exp\left(\frac{\hat{p}^{2}-m_{D^{*0}}^{2}}{\Lambda_{D^{*0}}^{2}}\right)\,$, (2.39)
$F^{(\gamma)}(q^{2})$ $=$ $\frac{\Lambda_{\gamma}^{2}}{\Lambda_{\gamma}^{2}-q^{2}}\,,\qquad\Lambda_{\gamma}=1\;{\rm GeV}\,.$ (2.40)
Here $\hat{p}^{2}<0$ and $q^{2}<0$. We take $\Lambda_{D^{*0}}=3.3-3.5\;{\rm GeV}$ estimated from a comparison of the model results to the experimental data (see Figs. 35).
The coupling constant $g_{D^{*0}D^{0}\gamma}$ is related to the decay width of $D^{*}(2007)^{0}\to D^{0}\gamma$ as follows
$\Gamma(D^{*0}\to D^{0}\gamma)=\frac{\alpha_{\rm em}}{24}\frac{(m_{D^{*0}}^{2}-m_{D^{0}}^{2})^{3}}{m_{D^{*0}}^{5}}|g_{D^{*0}D^{0}\gamma}|^{2}\,.$ (2.41)
From [1] we find
$m_{D^{*0}}=2006.85\pm 0.05~{\rm MeV}\,$,
$m_{D^{0}}=1864.84\pm 0.05~{\rm MeV}\,$,
$\Gamma(D^{*0}\to D^{0}\gamma)/\Gamma_{D^{*0}}=(35.3\pm 0.9)\%\,$,
$\Gamma_{D^{*0}}<2.1~{\rm MeV}\,.$ (2.42)
There is no accurate experimental result for the total width of $D^{*}(2007)^{0}$. However, there are some estimates, see e.g. [41, 42, 43, 44]. With $\Gamma_{D^{*0}}=55.3$ keV [43, 45] we get $|g_{D^{*0}D^{0}\gamma}|=5.97$ and this value is used in our calculations. The sign of $g_{D^{*0}D^{0}\gamma}$ does not matter as our amplitude is quadratic in the coupling.
We use in (2.36) $\tilde{\Delta}^{(D^{*0})\,\kappa_{1}\kappa_{2}}(s_{34},\hat{p}^{2})$ the propagators for the reggeized vector meson $D^{*}(2007)^{0}$. Two scenarios for the vector-meson reggeization effect were considered in [46, 47]. In the present work, we proceed analogously to (3.20)–(3.22) of [46]. The $D^{*0}$-meson propagator and the transverse function reads
$\tilde{\Delta}^{(D^{*0})\,\kappa_{1}\kappa_{2}}(s_{34},\hat{p}^{2})=-g^{\kappa_{1}\kappa_{2}}\tilde{\Delta}_{T}^{(D^{*0})}(s_{34},\hat{p}^{2})\,$, (2.43)
$\tilde{\Delta}_{T}^{(D^{*0})}(s_{34},\hat{p}^{2})=\Delta_{T}^{(D^{*0})}(\hat{p}^{2})\left(\exp(i\phi(s_{34}))\,\frac{s_{34}}{s_{\rm thr}}\right)^{\alpha_{D^{*}}(\hat{p}^{2})-1}\,$, (2.44)
where we take the simple expression $(\Delta_{T}^{(D^{*0})}(\hat{p}^{2}))^{-1}=\hat{p}^{2}-m_{D^{*0}}^{2}$, and
$\phi(s_{34})=\frac{\pi}{2}\exp\left(\frac{s_{\rm thr}-s_{34}}{s_{{\rm thr}}}\right)-\frac{\pi}{2}\,,\qquad s_{\rm thr}=4m_{D^{0}}^{2}\,.$ (2.45)
We use a specific nonlinear Regge trajectory for the $D^{*0}$ mesons, the so-called “square-root” trajectory [48], viz.
$\alpha_{D^{*}}(\hat{p}^{2})=\alpha_{D^{*}}(0)+\gamma\left(\sqrt{T_{D^{*}}}-\sqrt{T_{D^{*}}-\hat{p}^{2}}\right)\,$, (2.46)
where $\gamma$ governs the slope of the trajectory and $T_{D^{*}}$ denotes the trajectory termination point. The parameters are fixed to be $\alpha_{D^{*}}(0)=-1.02$, $\gamma=3.65$ GeV-1, $\sqrt{T_{D^{*}}}=3.91$ GeV; see Table I of [48].
II.4 Continuum production of charged $D$ mesons: $\gamma^{*}\gamma^{*}\to D^{+}D^{-}$
The continuum amplitude with the $D^{\pm}$ exchanges can be written in a similar manner as described in [49, 50] for the $\gamma\gamma\to H^{+}H^{-}$ and $\gamma\gamma\to p\bar{p}$ collisions, respectively. The amplitude for the subprocess $\gamma\gamma\to D^{+}D^{-}$ has the form:
${\cal M}_{\mu\nu}(\gamma^{*}\gamma^{*}\to D^{+}D^{-})$ $=$ $-ie^{2}\left[(q_{1}-2p_{3})_{\mu}(q_{1}-p_{3}+p_{4})_{\nu}\frac{1}{\hat{p}_{t}^{2}-m_{D}^{2}}+(q_{1}-2p_{4})_{\mu}(q_{1}+p_{3}-p_{4})_{\nu}\frac{1}{\hat{p}_{u}^{2}-m_{D}^{2}}-2g_{\mu\nu}\right]$ (2.47)
$\times F^{(\gamma)}(q_{1}^{2})F^{(\gamma)}(q_{2}^{2})F(\hat{p}_{t}^{2},\hat{p}_{u}^{2},p_{34}^{2})\,.$
Here $F^{(\gamma)}(q_{1,2}^{2})$ are given by (2.40) and the off-shell meson dependences are taken into account via multiplication of the amplitude by a common form factor $F(\hat{p}_{t}^{2},\hat{p}_{u}^{2},p_{34}^{2})$ [see (2.12), (2.13) of [50] but with the replacements $m_{p}\to m_{D^{+}}$ and $\Lambda_{p}\to\Lambda_{D}$]. The parameter $\Lambda_{D}$ should be fitted to the experimental data. In the present paper, we performed calculations for $F(\hat{p}_{t}^{2},\hat{p}_{u}^{2},p_{34}^{2})=1$.
By a similar way as in Sec. II.3, the $D^{+}D^{-}$-production mechanism via the $D^{*}(2010)^{\pm}$ exchanges can also be considered. To estimate the coupling constant $g_{D^{*}(2010)^{+}D^{+}\gamma}$ we find [1]
$m_{D^{*}(2010)^{+}}=2010.26\pm 0.05~{\rm MeV}\,$,
$m_{D^{+}}=1869.66\pm 0.05~{\rm MeV}\,$,
$\Gamma(D^{*}(2010)^{+}\to D^{+}\gamma)/\Gamma_{D^{*0}(2010)^{+}}=(1.6\pm 0.4)\%\,$,
$\Gamma_{D^{*}(2010)^{+}}=83.4\pm 1.8~{\rm keV}\,.$ (2.48)
Using (2.41) with (2.48) we get $|g_{D^{*}(2010)^{+}D^{+}\gamma}|=0.3$. Thus, the corresponding contribution of the continuum for $e^{+}e^{-}\to e^{+}e^{-}D^{+}D^{-}$ via the $D^{*}(2010)^{\pm}$ exchanges is much smaller than for the $D^{0}\bar{D}^{0}$ case, and can be safely neglected.
III Results and discussion
The experimental cross section (efficiency corrected) for $\chi_{c2}(3930)$ production, determined by the BaBar Collaboration (Eq. (17) of [5]), is
$\sigma_{\rm exp}(e^{+}e^{-}\to e^{+}e^{-}(\gamma\gamma\to\chi_{c2}(3930)\to D\bar{D}))=741\pm 166\;{\rm fb}\,.$ (3.1)
Here the error is only statistical. It is not clear whether the BaBar Collaboration imposed the cut on $p_{t}$ of the $D\bar{D}$ pair when presenting the integrated cross section.
The BaBar [5] and Belle [4] Collaborations conclude that
[…]
with the assumption for the Belle result that $B(\chi_{c2}\to D^{+}D^{-})=0.89\times B(\chi_{c2}\to D^{0}\bar{D}^{0})$. The PDG average gives $\Gamma_{\gamma\gamma}\times B(\chi_{c2}\to D\bar{D})=0.21\pm 0.04$ keV [1].
III.1 $e^{+}e^{-}\to e^{+}e^{-}\chi_{c2}(3930)$
We made calculations of the cross sections for the $e^{+}e^{-}\to e^{+}e^{-}\chi_{c2}(3930)$ reaction for c.m. energy $\sqrt{s}=10.54$ GeV. The results are $\sigma(e^{+}e^{-}\to e^{+}e^{-}\chi_{c2})=1.577$ pb using the Cornell potential, and $1.022$ pb using Buchmüller-Tye (BT) potential. To obtain (3.1), we find $B(\chi_{c2}\to D\bar{D})$ corresponding to $\Gamma_{\gamma\gamma}$ from Table 1 for these two potentials. We got
$\Gamma_{\gamma\gamma}\times B(\chi_{c2}\to D\bar{D})=\begin{cases}0.840~{\rm keV}\times 0.47^{+0.10}_{-0.11}=0.39^{+0.09}_{-0.09}~{\rm keV}&{\rm for~Cornell}\,,\\ 0.544~{\rm keV}\times 0.73^{+0.16}_{-0.17}=0.40^{+0.08}_{-0.10}~{\rm keV}&{\rm for~BT}\,.\end{cases}$ (3.3)
Comparing the results (3.3) to the BaBar result (3.2) ($\Gamma_{\gamma\gamma}\times B(\chi_{c2}(2P)\to D\bar{D})=0.24\pm 0.06~{\rm keV}$) one can see that they are roughly consistent within the uncertainty.
Taking into account the condition $p_{t,\chi_{c2}}<0.05$ GeV we obtain $\sigma(e^{+}e^{-}\to e^{+}e^{-}\chi_{c2}(2P))=0.625$ pb (Cornell) and $0.405$ pb (BT). These results with comparison to (3.2) leads to $B(\chi_{c2}\to D\bar{D})=1.19$ and 1.83 for Cornell and BT potential, respectively. This suggests that the value given in (3.1) does not include cuts.
III.2 $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$
In Table 3, we listed the values of the fitted parameter $|g_{\chi_{c2}D\bar{D}}|$ obtained from comparison to the combined BaBar cross section (3.1). For neutral and charged channel we obtain
$\sigma(e^{+}e^{-}\to e^{+}e^{-}(\chi_{c2}(3930)\to D^{0}\bar{D}^{0}))$ $=390\pm 87\;{\rm fb}\,$,
$\sigma(e^{+}e^{-}\to e^{+}e^{-}(\chi_{c2}(3930)\to D^{+}D^{-}))$ $=351\pm 79\;{\rm fb}\,.$ (3.4)
The requirement of very small transverse momentum $p_{t,D\bar{D}}<0.05$ GeV is motivated by an untagged analysis of $\gamma\gamma\to D\bar{D}$ production by Belle and BaBar. Due to the $p_{t,D\bar{D}}$ cut the cross section is reduced by a factor of 2.54. In order to describe the BaBar result (3.1), this causes a larger value for the constant $|g_{\chi_{c2}D\bar{D}}|$ and thus leads to branching fraction $B(\chi_{c2}\to D\bar{D})\sim 1$. Then, the theoretical results for the potential models under consideration are not consistent with the experiment.
Table 3: Results for the reaction $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$ via $\chi_{c2}(3930)$ calculated for $\sqrt{s}=10.54$ GeV without and with the limitations on the transverse momenta of the final state particles. The fitted coupling parameter $|g_{\chi_{c2}D\bar{D}}|$ for the Cornell and Buchmüller-Tye (BT) potential is shown. In the last column the product of the two-photon width of the $\chi_{c2}(3930)$ state and the branching fraction to $D\bar{D}$ for each model is given.
Potential $|g_{\chi_{c2}D\bar{D}}|$ $B(\chi_{c2}\to D\bar{D})$ $\Gamma_{\gamma\gamma}\times B(\chi_{c2}\to D\bar{D})$ (keV)
no cuts Cornell $8.1^{+0.9}_{-1.0}$ $0.38\pm 0.08$ $0.32\pm 0.07$
BT $10.1^{+1.1}_{-1.2}$ $0.58\pm 0.13$
$p_{t1},p_{t2}<1$ GeV Cornell $8.4^{+0.9}_{-1.0}$ $0.40\pm 0.09$ $0.34\pm 0.08$
BT $10.4^{+1.1}_{-1.2}$ $0.62\pm 0.14$
$p_{t,D\bar{D}}<0.05$ GeV Cornell $12.9^{+1.4}_{-1.5}$ $0.96\pm 0.21$ $0.80\pm 0.18$
BT $16.1^{+1.7}_{-1.9}$ $1.48\pm 0.33$
From Table 3 one can see that the obtained value of the branching fraction $B(\chi_{c2}(2P)\to D\bar{D})$ depends on the potential model (the two-photon width of the $\chi_{c2}(3930)$ state) assumed in the calculations. For the results calculated without kinematical cuts, the obtained value for the product $\Gamma_{\gamma\gamma}\times B(\chi_{c2}(2P)\to D\bar{D})$ is consistent with existing Belle and BaBar measurements within uncertainties. From our analysis, the corresponding branching fraction is $B(\chi_{c2}(3930)\to D\bar{D})=0.38\pm 0.08$ for the Cornell potential and $0.58\pm 0.13$ for the Buchmüller-Tye potential. The latter value is compatible with the one reported in [51].
Now we wish to show some differential distributions for the $e^{+}e^{-}\to e^{+}e^{-}(\chi_{c2}(3930)\to D^{0}{\bar{D}}^{0})$ reaction. Figure 2 shows the distributions in $p_{t,D\bar{D}}$ the transverse momentum of the $D^{0}{\bar{D}}^{0}$ pair, in $\log_{10}(p_{t1}/1\,\rm{GeV})$ one of the integration variables, and in $|t_{1}|$ the four-momentum transfer squared. In the calculation we used $g_{\chi_{c2}D\bar{D}}=10.1$ corresponding to the BT potential. We show results when we have not imposed any restrictions on kinematic variables and the results with the condition $p_{t,D\bar{D}}<0.05$ GeV. The distribution in $p_{t,D\bar{D}}$ peaks in the region of low $p_{t,D\bar{D}}$. The $t$ distributions are strongly peaked at very small $|t_{1}|$ and $|t_{2}|$. This is caused by the factors $1/t_{1,2}$ from the photon propagators. One can clearly see that imposing the condition on $p_{t,D\bar{D}}<0.05$ GeV significantly reduces the cross-section and limits the momentum transfers to the electrons, $|t_{1,2}|<0.01$ GeV2.
Figure 2: Distributions in $p_{t,D\bar{D}}$, $\log_{10}(p_{t1}/1\,\rm{GeV})$ and $|t_{1}|$ for the $e^{+}e^{-}\to e^{+}e^{-}(\chi_{c2}(3930)\to D^{0}{\bar{D}}^{0})$ reaction for $\sqrt{s}=10.54$ GeV. We show the results without and with the cut on $p_{t,{\rm sum}}<0.05$ GeV. The calculations were done for the BT potential and with $g_{\chi_{c2}D\bar{D}}=10.1$.
We compare our theoretical results for the $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$ reaction calculated for $\sqrt{s}=10.54$ GeV with the Belle data from Fig. 2 of [4] and BaBar data from Fig. 5 of [5]. The $\chi_{c0}(3860)$ and $\chi_{c2}(3930)$ contributions are calculated for the BT potential with $R^{\prime}_{2P}(0)=0.326$ GeV5/2 (see Tables 1 and 2) and with $g_{\chi_{c2}D\bar{D}}=10.1$ (see Table 3), and $g_{\chi_{c0}D\bar{D}}=2.52$. In Fig. 3(a)-(c) we show the invariant mass distributions for the combined $D\bar{D}$ final state and for the $D^{0}\bar{D}^{0}$ and $D^{+}D^{-}$ channels separately. We can see that the complete model results, containing the continuum and resonant contributions, can reasonably well describe the experimental data. There is a significant interference effect between the $\chi_{c0}(3860)$ resonance and continuum terms. The $D^{0}\bar{D}^{0}$-continuum contribution increases from the reaction threshold, reaches a maximum at $M_{D^{0}\bar{D}^{0}}\sim 3.8$ GeV, and has a large share in the $\chi_{c2}(3930)$ resonance region. In these calculations, for the $D^{0}\bar{D}^{0}$-continuum contribution, we use $\Lambda_{D^{*0}}=3.5~{\rm GeV}$ in (2.39). From the presented results, it is clearly visible that the $D^{+}D^{-}$-continuum contribution is small. Since the form factor $F(\hat{p}_{t}^{2},\hat{p}_{u}^{2},p_{34}^{2})$ is set to 1 in the calculation, our estimate should be viewed as an upper limit, especially at large $M_{D^{+}D^{-}}$. In Fig. 3(d) we show the distribution in $p_{t,D\bar{D}}$ in the region $M_{D\bar{D}}\in(3.91,3.95)$ GeV.
(a)Refer to caption (b)Refer to caption (c)Refer to caption (d)Refer to caption
Figure 3: The invariant mass distribution for the $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$ reaction for the neutral and charged channels, and the combined $D\bar{D}$ final state. (a) Comparison of theoretical results to the Belle data [4]. (b) Comparison of theoretical results to the BaBar data [5]. (c) The differential cross sections $d\sigma/dM_{D\bar{D}}$ for the complete model (continuum, $\chi_{c0}(3860)$, and $\chi_{c2}(3930)$ contributions) and for the continuum contribution alone. For the $\chi_{c0}(3860)$ term we take (2.25). The red long-dashed lines correspond to the $D^{0}{\bar{D}}^{0}$ channel and the blue dashed lines correspond to the $D^{+}D^{-}$ channel. The black solid line represents the sum of all contributions for two channels. In panel (d), we show the differential distributions in transverse momentum of the $D\bar{D}$ pair. The calculations were done for $\sqrt{s}=10.54$ GeV and with the cut on $M_{D\bar{D}}\in(3.91,3.95)$ GeV. The meaning of the lines is the same as in Fig. 3(c).
For illustrative purposes, in Fig. 4 we present our results obtained for $\Gamma_{\chi_{c0}}=200$ MeV (2.24) for the $\chi_{c0}(3860)$ resonance. In this case, we have that the increase in $\Gamma_{\chi_{c0}}$ implies a worse description of low-$M_{D\bar{D}}$ data (especially for the $D^{+}D^{-}$ BaBar data). Here, we keep the coupling constants the same as in the previous scenario. In general, using the larger total width, the coupling constant of $\chi_{c0}\to D\bar{D}$ could be consistently re-estimated. If a larger value for $B(\chi_{c0}\to D\bar{D})$ is used, this can be compensated for by a smaller two-photon width (see Table 2). From the right panel of Fig. 4 we see that the contributions from the $\chi_{c2}(3930)$ resonance and $D^{0}\bar{D}^{0}$-continuum are dominant. The term from $\chi_{c0}(3860)$ is found to be small but the interference effect between the $\chi_{c0}(3860)$ and the $D^{0}\bar{D}^{0}$ continuum is significant.
Figure 4: Comparison of theoretical results to the BaBar data [5]. The meaning of the lines is as in Fig. 3, but for the $\chi_{c0}(3860)$ term we used the middle values from (2.24).
An additional comment is in order here. We are not sure whether the Belle data shown in Fig. 3 was corrected for experimental efficiencies. If not, this could partly explain why our results are higher than the Belle data. Therefore, we have also compared our model results to the efficiency-corrected BaBar data, which was provided only for the combined $D\bar{D}$ final state.
Finally, in Fig. 5 we show our complete model results including $D^{0}\bar{D}^{0}$ continuum, $D^{+}D^{-}$ continuum, $\chi_{c0}(3860)$, and $\chi_{c2}(3930)$ contributions together with the efficiency-corrected (combined) BaBar data taken from Fig. 10 of [5]. In the calculation we used the Buchmüller-Tye potential for the excited $2P$ states, $\chi_{c0,2}(2P)$. The $D^{0}\bar{D}^{0}$-continuum contribution from $t/u$-channel vector-meson $D^{*}(2007)^{0}$ exchanges depends on the parameter of the off-shell form-factor (2.39). In the calculations we take two values of the form-factor parameter $\Lambda_{D^{*0}}=3.3~{\rm GeV}$ and $3.5~{\rm GeV}$ corresponding to the lower and upper line in the bands shown in Fig. 5, respectively. The presented calculations correspond to two scenarios of the $\chi_{c0}(3860)$ total width: (1) $\Gamma_{\chi_{c0}}=50$ MeV with $\chi^{2}/{\rm dof}=2.44$ (solid upper line) and $\chi^{2}/{\rm dof}=1.61$ (solid lower line); (2) $\Gamma_{\chi_{c0}}=200$ MeV with $\chi^{2}/{\rm dof}=2.61$ (dashed upper line) and $\chi^{2}/{\rm dof}=2.06$ (dashed lower line). We also checked that limiting to the range of $2m_{D}\leq M_{D\bar{D}}<3.9~{\rm GeV}$ the resulting values of $\chi^{2}/{\rm dof}$ are 1.82–2.04 and 1.39–2.85 for scenarios (1) and (2), respectively.
Figure 5: Comparison of our complete model results to the efficiency-corrected (combined) BaBar data from Fig. 10 of [5]. The black solid lines represent results obtained for $\Gamma_{\chi_{c0}}=50$ MeV (2.25), while the blue dashed lines correspond to $\Gamma_{\chi_{c0}}=200$ MeV (2.24). Here, the calculations for the $D^{0}\bar{D}^{0}$-continuum contribution were made for $\Lambda_{D^{*0}}=3.3$ and $3.5~{\rm GeV}$ in (2.39), which correspond to the lower and upper lines in the bands, respectively. We used the Buchmüller-Tye potential for the excited $2P$ states, $\chi_{c0,2}(2P)$.
IV Conclusions
In the present paper, we have concentrated on the possible observation of excited $P$-wave charmonia in the $e^{+}e^{-}\to e^{+}e^{-}D\bar{D}$ reaction. One of our goals was to estimate the cross-section of meson-pair continua for both $D^{0}{\bar{D}}^{0}$ and $D^{+}D^{-}$ channels separately. We have considered $t$- and $u$-channel exchanges of the vector $D^{*0}$ and $D^{*\pm}$ mesons. The corresponding coupling constants for the $D^{*}D\gamma$ couplings have been obtained from the experimental radiative partial decay widths of $D^{*}\to D\gamma$. For the $D^{+}D^{-}$ channel, we have also considered the continuum mechanism via the $D^{\pm}$ $t/u$-channel exchanges plus contact term. Comparing our model results with the experimental data (Belle [4], BaBar [5]) we find a large contribution of the continuum to the $D^{0}{\bar{D}}^{0}$ channel and a much smaller one for the $D^{+}D^{-}$ channel (see Fig. 3).
Our analysis suggests that the broad bump at $M_{D\bar{D}}=3.8$ GeV observed in the $D^{0}{\bar{D}}^{0}$ channel, and not observed significantly in $D^{+}D^{-}$, is rather due to the continuum mechanism. However, a certain contribution of a broad resonance $\chi_{c0}(3860)$ cannot be completely ruled out. We have shown that taking this resonance into account with $\Gamma_{\chi_{c0}}=50$ MeV improves the description of the near-threshold data, in particular, the BaBar data for the $D^{+}D^{-}$ channel shown in Fig. 3(b). It is rather difficult to understand a large difference between $\Gamma_{\chi_{c0}}$ and $\Gamma(\chi_{c0}\to D\bar{D})$ in the scenario with large $\Gamma_{\chi_{c0}}$ given in (2.24) which leads to surprisingly small branching fraction $B(\chi_{c0}\to D\bar{D})$. It should be noted that the potential-model predictions for the mass and width of the $\chi_{c0}(2P)$ state differ from the PDG averages. This indicates that further theoretical and experimental research is needed to understand properties of the $\chi_{c0}(2P)$ state more precisely.
We have also estimated the cross section for the $\chi_{c2}(3930)$ production in no tag conditions using $R_{2P}^{\prime}(0)$ calculated within the light front quark model. Calculations were performed for different potentials known from the literature; see Table 1. We have obtained an integrated cross-section $\sigma(e^{+}e^{-}\to e^{+}e^{-}\chi_{c2}(3930))\sim 1$ pb using the Buchmüller-Tye potential. From a comparison of our results for the $e^{+}e^{-}\to e^{+}e^{-}(\chi_{c2}(3930)\to D\bar{D})$ reaction to the BaBar results, we have found the branching fraction $B(\chi_{c2}(3930)\to D\bar{D})=0.58\pm 0.13$ (Buchmüller-Tye potential) and $0.38\pm 0.08$ (Cornell potential). We have also obtained $\Gamma_{\gamma\gamma}\times B(\chi_{c2}\to D\bar{D})=0.32\pm 0.07$ keV (see Table 3). This is not far from the experimental value (3.2).
Our results strongly motivate a more detailed analysis of the $D\bar{D}$ production. This would require Monte Carlo simulations of both neutral and charged channels, including experimental acceptances and efficiencies. Hopefully, this will be possible with future Belle II high-statistics data.